OpenStudy (anonymous):

MEDAL WILL BE AWARDED AND FAN! Determine the type of boundary line and shading for the graph of the inequality 5x + y greater than or equal to -10. Dashed line with shading on the side that includes the origin Solid line with shading on the side that does not include the origin Dashed line with shading on the side that does not include the origin Solid line with shading on the side that includes the origin

OpenStudy (anonymous):

First off, you know that it is greater than OR EQUAL TO. Does this mean it is a dashed or solid line?

OpenStudy (anonymous):

um, solid

OpenStudy (anonymous):

Correct, so can already rule out two options. Now you have\[5x +y \ge-10\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What you want to do is isolate y. To do this, subtract 5x from both sides.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wouldn't we just have 5xy now?

OpenStudy (anonymous):

Can you show me what you have after you subtract 5x from both sides?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

i know 5x - 5x is zero so thats gone

OpenStudy (anonymous):

\[y \ge-10-5x\] this is what you have now.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

This is also\[y \ge-5x-10\] which is slope intercept form.

OpenStudy (anonymous):

oh, ok i see

OpenStudy (anonymous):

You shade above the line because the sign is a greater than. Since you shade above the line, this includes the origin. Does this help?

OpenStudy (anonymous):

so its b?

OpenStudy (anonymous):

no no no

OpenStudy (anonymous):

d sorry

OpenStudy (anonymous):

Yes! :D

OpenStudy (anonymous):

thank you so much dude, im gonna give you a medal and fan you

OpenStudy (anonymous):

Thank u! :)

OpenStudy (anonymous):

im going to close this question now

OpenStudy (anonymous):

K. Anytime you need help with something else, just message me. :)

OpenStudy (anonymous):

ok